OpenStudy (hulkart):

Evaluate the logarithm with a calculator: log2(20). 2.9957 1.3010 4.3219 0.3010

Nnesha (nnesha):

change of base formula \[\huge\rm log_a b = \frac{ \log b }{ \log a }\]

OpenStudy (hulkart):

4.3219?

Nnesha (nnesha):

yep

OpenStudy (hulkart):

how do you get it tho

OpenStudy (hulkart):

?

Nnesha (nnesha):

how did YOU get it ?

OpenStudy (hulkart):

math.

OpenStudy (hulkart):

:D

Nnesha (nnesha):

apply the change of base formula and then use calculator

OpenStudy (hulkart):

yeah...what I did:D

Nnesha (nnesha):

so yeah ....

OpenStudy (hulkart):

what? lol